OpenStudy (anonymous):

What is the factored form of the expression? 12d2 + 4d – 1

OpenStudy (callisto):

12d^2 + 4d – 1 =12d^2 + 6d - 2d - 1 = 6d(2d+1) - (2d+1) = ...?

OpenStudy (anonymous):

thanks